mikewwe13:

The perimeter of Dana's rectangular garden is 61 feet. The length of her garden is 5 feet more than twice the width Create an equation or equations that can be used to determine the length and width of the garden. Be sure to define your variables. What is the length and width of the garden? Show all your work. Enter your equation(s), definition of variables, answers, and work in the box provided.

Vocaloid:

if we let L = length and W = width how do you think we can write: "The length of her garden is 5 feet more than twice the width" in terms of variables?

mikewwe13:

yes

Vocaloid:

I am not asking a yes or no question, I am asking for your opinion

mikewwe13:

well for my opinion i think it should be just that, the best way to write it

Vocaloid:

Again, I am asking, how can we write ""The length of her garden is 5 feet more than twice the width" in terms of variables?" using our symbols L and W ?

Vocaloid:

how can we write "twice the width" using numbers and variables instead of words?

mikewwe13:

to write twice the width is by saying the length of a rectangle is twice the width. so therefore The number of square units in it's area is four times the number of units in it's perimeter

Vocaloid:

ok, but how would we write that using Math, not words?

Vocaloid:

hint: "twice the width" means multiply width by 2 how can we write this using VARIABLES and NUMBERS not words?

Vocaloid:

if W = width, then what is "twice the width"?

mikewwe13:

The length of a rectangle

Vocaloid:

hint: "twice" means "multiply by 2"

Vocaloid:

how can we say "twice the width" using MATH not english?

mikewwe13:

by multiplying by 2

Vocaloid:

if W is width, then how can we express "multiplying width by 2" WITHOUT using the words "multiply" or "width"

mikewwe13:

area of a rectangle

Vocaloid:

hint: 2 multiplied by x is 2x 2 multiplied by W is...?

mikewwe13:

2x

Vocaloid:

no

Vocaloid:

2 multiplied by W is...?

mikewwe13:

2

Vocaloid:

follow the pattern. Think slowly and carefully before you answer. 2 multiplied by A is 2A 2 multiplied by T is 2T 2 multiplied by smiley face is 2smileyface 2 multiplied by W is ___ fill in the blank.

mikewwe13:

twice the width" means multiply width by 2

Vocaloid:

notice the first line? 2 multiplied by A is 2A notice how we took 2 and A and wrote them together to multiply them? 2A 2 multiplied by T is 2T notice how we took 2 and T and put them together to get 2T? so what do you think we should write if we want 2 multiplied by W?

mikewwe13:

2W

mikewwe13:

i'm sorry i'm causing you to be stressed but i just need help on my work

mikewwe13:

the answer is 2W

mikewwe13:

i didn't get it at first because my brain was at somewhere but i got it

Vocaloid:

Ok, I'm back

Vocaloid:

It says that the garden is "five feet more than twice the length" So if width = 2W how can we express length in terms of math? Not words, just variables and numbers

Vocaloid:

Hint: "5 more than" implies + 5 (Addition)

mikewwe13:

w + h = (w+h) - h = - h w = - h w = w = =

Vocaloid:

Where are you getting h? We only have two variables, w and L

mikewwe13:

i mean with l but it was malfunction with the keyboard

Vocaloid:

"Five more than twice the width" means "5 + 2W"

Vocaloid:

Now we use W and L to write an expression for perimeter

Vocaloid:

Perimeter = 2L + 2W Since L = 2W + 5 we can replace L with "2W + 5" into P

Vocaloid:

Perimeter = 2(W+5) + 2W = 61 Solve for W

mikewwe13:

that's it ?

Vocaloid:

You need to solve that for W

mikewwe13:

w = 51/4

Vocaloid:

Oh whoops made a typo

Vocaloid:

2(2W+ 5) + 2W = 61

Vocaloid:

Solve that for W

mikewwe13:

W = 17/2

Vocaloid:

Good now find L if L = 2W + 5

mikewwe13:

l=2w+5

Vocaloid:

Take the value you got for W and substitute that into 2W + 5

Vocaloid:

If W = 17/2 then 2W + 5 =?

mikewwe13:

the answers for substituting them is l = 22 and w = 17/2

mikewwe13:

second question is 2w + 5

Vocaloid:

Sure

mikewwe13:

what else ?

Vocaloid:

That's it

Vocaloid:

Just write down what we have done so far and organize the work